D Question 1 3 pts Consider a firm with the following parameters: Physical Lifespan Capital Expense, Year 0 EBITDA/Year Overall Firm Cost of Capital Tax Rate Loan Interest Payments Depreciation Project Life 7 years $13,000 $9,000 9AY。 38% $16,000 $551 4 years What is the total PV of this firm, including the loan? Question 2 3 pts Your company purchases two machines, which cost $3,000 and $6,000, respectively. They have inear straight-line depreciation schedules of 7 and 4 years, respectively. What is the present value of the depreciation, given a cost of capital of 6.7%?Please answer both. Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Answer 1

1 2 3 4 5 6 7
EBITDA 9000 9000 9000 9000 9000 9000 9000
(-): Depreciation 3250 3250 3250 3250
(-): Interest Expense 551 551 551 551 551 551 551
EBT 5199 5199 5199 5199 8449 8449 8449
(-): Tax 1975.62 1975.62 1975.62 1975.62 3210.62 3210.62 3210.62
Earning 3223.38 3223.38 3223.38 3223.38 5238.38 5238.38 5238.38
Add: Depreciation 3250 3250 3250 3250
Cash Flow 6473.38 6473.38 6473.38 6473.38 5238.38 5238.38 5238.38

PV of firm = -13000+16000+6473.38*.911+6473.38*.829+6473.38*0.755+6473.38*0.688+5238.38*0.627+5238.38*0.571+5238.38*0.519-16000*0.519

=3000+20604.7685+7686.5945-8304

=$22987.363

Answer 2

Present Value of Depreciation

=428.5714*5.44461+1500*3.4102

=2334.0427+5115.30

=7449.3427

Note : Depreciation = Asset Cost/ Life of asset